### v1

@@ III-A Feed-Forward - Humanoid
As stated before, the first experiment serves as a replication experiment for the original OpenAI-ES paper [8], as well as a sanity check of our implementation. It consists of training a feed-forward model using OpenAI-ES in Humanoid environment. We can see in Figure 2 that our implementation is functional and is comparable in performance with the original implementation. The progression of the training shown in Figure 2(a) also serves as a benchmark against which the results of other experiments can be compared.
@@ III-B Decision Transformer - Humanoid
In this experiment, we inspected the ability of the given evolution strategy to train even the more complicated model, which the Decision Transformers are. We trained a Decision Transformer model using OpenAI-ES, again in the Humanoid environment. Because the model is almost five times larger, we quadrupled the size of the population the algorithm works with. The results are shown in Figure 3(a).
We also explored whether it is necessary to use such a substantial population or whether half of it would be sufficient. The results of this ablation study can be seen in Figure 3(b). We can see that even though the algorithm is sometimes still capable of training the model, more often than not it does not succeed in doing so.
(a) Evaluation results
(b) Evaluation wall-clock times
Figure 5: OpenAI-ES used on a Decision Transformer model for Atari game Hero in ten runs of the experiment. Figure 5(a) shows us the evaluation results after every iteration of the algorithm. In Figure 5(b), we can see the wall-clock time the algorithm needed each iteration in seconds.
@@ III-C Decision Transformer - Humanoid - Pretrained
Yet, this approach has its costs. First, the OpenAI-ES uses virtual batch normalization (VBN) [19]. This changes the inputs to the model as the training progresses. So, either we would have to use VBN even during the pretraining, or we cannot use the VBN during the following training with the evolution strategy. We opted for the second possibility, as it is more general and allows us to showcase a further training of any model we might have using the evolution strategy, even though the first option would be more favorable for the algorithm, as the VBN improves its reliability [8] – it should, however, be most important, when the model is mostly random at the beginning of the training, according to the original authors.
Second, using the same values for learning rate and noise deviation hyperparameters as when training from scratch leads to complete randomization of the model at the beginning of the training and then the training proceeds as if from scratch. So, for the pretraining to have an effect, we have to decrease both values (to 0.01 in our experiment). This, in turn, leads to a slower progression of the training.
Hence, again in the Humanoid environment, we pretrained the Decision Transformer using a behavior cloning to a feed-forward model trained using SAC algorithm [20], before training it further using OpenAI-ES. The feed-forward model and the pretrained Decision Transformer had a fitness value around 4. The results of the subsequent training of the pretrained Decision Transformer using the evolution strategy are shown in Figure 4(a).
Again, we explored the possibility of utilizing a smaller population for this training. As can be seen in Figure 4(b), this time, although the reliability of the training is again diminished, most of the time the algorithm works surprisingly well compared to how it works with the full population.
Next, we tried using standard (i.e. the same as in the previous experiment without the pretraining) values for learning rate and noise deviation hyperparameters both with and without the VBN. Without the VBN the training showed no progression at all, and with the VBN the training started to show similar patterns as the one from the previous experiment, yet still was maybe a bit slower. The corresponding plots are quite uninteresting, but together with the underlying experiment data may be found along the code and other data for this paper.
@@ III-D Decision Transformer - Atari
As a final experiment, we tested the ability of OpenAI-ES to train an even larger Decision Transformer in Atari games environment, more specifically, in the game Hero. Again, since the model is approximately three times bigger than in the previous experiments, we further doubled the population size for the algorithm. Even though Figure 5(a) shows that a gradual improvement of the model really does occur, Figure 5(b) reveals that it is accompanied by quite a significant increase in the time required to process a single iteration. This is due to slower model inference, as the model is substantially larger, as well as longer episodes in the environment.
@@ IV Discussion
As we can see by comparing Figures 2(a) and 3(a), the larger Decision Transformer model has an fitness advantage during the training over the smaller feed-forward model. But what is more important for this article, OpenAI-ES appears capable of training it without substantial problems, although the process takes longer. However, the training is less robust, as evidenced in Figure 3(a), which shows that during one experiment run, the model failed to achieve any significant improvement. For greater robustness, we would probably have to further increase the population size.
@@
Regarding our specific usage of the algorithm on Decision Transformers, we need to settle one thing. As we introduced in Section II-B, Decision Transformer operates with return-to-go tokens, which in the original paper [10] serve as a desired return the agent should achieve in a remainder of an episode. This works well when training in the original offline reinforcement learning manner, but when using the evolution strategy to train the Decision Transformers in our online setting, there is no pressure to take these tokens into account, and thus the tokens are then ignored. Nonetheless, we believe that only a minor modification of the algorithm may induce the agent to pay attention to these tokens. Let us propose an outline of such modifications for possible future work. Each evaluation would consist of several subevaluations; each subevaluation would be assigned a desired return sampled from a $\mathcal{N}(\mu,\sigma)$, a normal distribution centered at $\mu$ with variance $\sigma$. The value of $\mu$ should be computed from the returns obtained during the previous iteration in such a manner that it is within an interval between the best return obtained and the mean return obtained during the last iteration, so an improvement is gradually made. The variance $\sigma$ would be a hyperparameter, or it could possibly be a variation of the returns obtained during the last iteration. The fitness of an individual would not be the mean of its returns obtained, as is the case in OpenAI-ES, but rather a decreasing function of the absolute values of the differences between the return obtained and the desired return in each subevaluation; ergo, the larger the differences, the smaller the fitness. This might be, e.g., the negative of a weighted sum of the differences.
@@
Even when doing so, we can see in Figure 4 that the model first gets "broken" – its performance worsens during first few iterations – before it starts to improve again. Nevertheless, the improvement comes in terms of iterations much sooner than when no pretraing is utilized. Furthermore, we can observe by inspecting rollouts of both the initial and final models that those two show similar gaits. This similarity gets even more accentuated, when compared to the diversity in final behaviors that are yielded from the experiment in Section III-B, so when the algorithm is run without the pretraining. Therefore, we can clearly see that when initialized by the pretrained model the training clearly builds upon the seeded model, despite its initial deterioration.
@@
A reason for the above-mentioned deterioration at the beginning of the training from pretrained model, as well as the need for the aforementioned change of hyperparameters is, we believe, the following: Gradient training – of which the behavior cloning is a representant – follows a different overall goal than the evolution strategy. It strives for the best possible parameters of the model it trains with regard to its loss function. In our case of behavior cloning, this means that we want to get our model to behave exactly as the one we are cloning towards. Nonetheless, this tells us nothing about how the model behaves when we perform a slight perturbation of parameters. Hence, after the perturbation, the model is likely to no longer function properly – or at least as well – anymore. So we can say that the model is brittle in a sense that even a slight change of parameters is likely to somehow break the model. This is, however, in stark contrast to how the evolution strategy operates and what it tries to achieve. The goal of the evolution strategy is not only that its current model performs as well as possible, but that we get the best possible behavior in expectation when sampling model parameters from its current distribution. Thus, the models trained by the evolution strategy might be considered more robust – that might help, e.g., in areas where we train on precise hardware, but the model needs to be afterwards deployed on a cheaper hardware with less precision, where the weights of the model need to be rounded. Ergo, the algorithm needs to first, let us say, "robustify" the model and only then it can train it towards a better performance.
@@
Regardless, in the end, the results shown in Section III-B suggest that the pretraining is not really necessary. But then what explains why the problem stated at the beginning of Section III-C does not occur in practice? In Figure 2(b), we can see that the algorithm improves not only by shifting model weights towards the few better performing individuals of the population, but mostly by shifting the weights away from the worst performing individuals, as the majority of the population consists of worse individuals. This might indicate why the algorithm does not really struggle with training larger and more complex models, where it is even harder to stumble onto a working solution just by adding some noise to the model parameters.
Finally, let us just briefly comment on Section III-D. Yes, the OpenAI-ES is capable of training even such larger models, but it just takes too long. Of course, this could be easily solved by even more massive paralelization over even more processes and CPUs – although 300 workers is already quite a lot and we believe there is a limit on the scaling. But the question lingers: How fast would a gradient algorithm – even more so some Temporal-Difference based [21] (Actor-Critic like [22]) – be able to train these large models? And on top of that, gradient algorithms work best on GPUs, which one might also use to accelerate an inference in such large models. Meanwhile, evolution strategies work with a lot of CPUs instead, which are mostly better for environment simulations, but worse for the model inference, therefore we have a slower inference during rollouts – the slower the larger the model we have. So, in essence, where is the boundary, when the length of an episode in an environment and the size of a model favor gradients over evolution? (And yes, this question is – just as similar questions have always been in history of artificial intelligence and even more so of machine learning – deeply dependent on a current state of hardware available. But we believe the hardware cannot be improved indefinitely.) An indirect answer to this question might be utilization of hybrid algorithms that combine strengths of both approaches, the gradient algorithms, and the evolution [23].
@@ V Conclusion
We have investigated the ability of OpenAI-ES (and, in extension, of similar evolution strategies) to train larger and more complex models than the standard feed-forward ones used in the literature so far in reinforcement learning. We have also suggested a method to help this training by pretraining the model using behavior cloning towards some smaller, possibly weaker, yet easier to train model. However, the evolution strategy proved to be capable of training large models, whereas the pretraining showed multiple flaws during the experiments. Still, it helped us shed light onto the field of hybridizing gradients and evolution strategies, as we have discussed based on our experiments. It helped us understand why the sequential combination of first training using gradients and then proceeding to train using similar distribution-based evolution strategies will not work well. And it also served to support the claim that the distributional evolution strategies like OpenAI-ES yield robust models, which are less prone, e.g., to rounding the weights due to low-precision hardware they might be deployed on.

### v2

@@ III-A Feedforward - Humanoid
As stated before, the first experiment serves as a replication experiment for the original OpenAI-ES paper [8], as well as a sanity check of our implementation. It consists of training a feedforward model using OpenAI-ES in Humanoid environment. The hyperparameters are thus mostly directly copied from the original paper, with the exception of those affected by our implementation changes (e.g., a learning rate) which were set manually through experiments. We can see in Figure 2 that our implementation is functional and is comparable in performance with the original implementation. The progression of the training shown in the figure also serves as a benchmark against which the results of other experiments can be compared.
TABLE I: Median and quartiles of the average (unrescaled) returns of the trained models with different initial return-to-go argument values.
Return-to-go Q1 return Median return Q3 return
1000-1000- 1000 7466 8447 8721
0 7467 8440 8730
7000 7505 8370 8672
1 000 000 7489 8354 8667
@@ III-B Decision Transformer - Humanoid
(a) Without pretraining
(b) With pretraining
Figure 3: Results of the ablation studies for training of Decision Transformers using OpenAI-ES done in the Humanoid environment for ten runs each, aggregated into medians and quartiles. Figure 3(a) shows the fitness progression of the best yet found model for a training without first pretraining the model, as introduced in Section III-B. “Ours” stands for our setting of hyperparameters; “Half population” is the same except the size of the population is cut in half. Figure 3(b) shows the best-yet fitness values for a training utilizing pretraining, as introduced in Section III-C. “Ours” stands for our setting of hyperparameters; “Half population” has the size of the population cut in half; “Standard, no VBN” has the same hyperparameters as when no pretraining was used, so the same as experiment in Section III-B, but the virtual batch normalization is not used; “Standard, with VBN” has exactly the same hyperparameters as when training without the pretraining.
In this experiment, we inspected the ability of the given evolution strategy to train even a more complicated model, which the Decision Transformers are. We trained a Decision Transformer model using OpenAI-ES, again in the Humanoid environment. Because the model is almost five times larger, we quadrupled the size of the population the algorithm works with. The results are shown in Figure 2. There, we can see that the algorithm is fully capable of training our larger model, achieving very good results using a constant number of iterations (albeit at the expense of a longer wall-clock runtime).
Because a larger population increases the time needed for the computation with the same available resources, a question arises whether such increase in population is truly necessary. Therefore, we also explored whether it is necessary to use such a substantial population or whether half of it would be sufficient. The results of this ablation study can be seen in Figure 3(a). We can see that even though the algorithm is sometimes still capable of training the model, more often than not it does not succeed in doing so.
Last, we want to check whether the trained models respond to different returns-to-go passed to the model at the beginning of each episode as the desired performance, the target return. Hence, we simulated the trained model in the environment with various intitial return-to-go values. The results of these experiments can be found in Table I. We can see there is no significant difference between the results of the models with distinct returns-to-go passed.
@@ III-C Decision Transformer - Humanoid - Pretrained
Yet, this approach has its costs. First, the OpenAI-ES uses virtual batch normalization (VBN) [27]. This changes the inputs to the model as the training progresses. So, either we would have to use VBN even during the pretraining, or we cannot use the VBN during the following training with the evolution strategy. We opted for the second possibility, as it is more general and allows us to showcase a further training of any model we might have using the evolution strategy, even though the first option would be more favorable for the algorithm, as the VBN improves its reliability [8] – it should, however, be reportedly most important when the model is mostly random at the beginning of the training.
Second, using the same values for learning rate and noise deviation hyperparameters as when training from scratch leads to a complete randomization of the model at the beginning of the training and then the training proceeds as if from scratch. So, for the pretraining to have an effect, we have to decrease both values (to 0.01 in our experiment). This, in turn, leads to a slower progression of the training.
Hence, again in the Humanoid environment, we pretrained the Decision Transformer using a behavior cloning to a feedforward model trained using SAC algorithm [28], before training it further using OpenAI-ES. The feedforward model and the pretrained Decision Transformer had a fitness value around 4. The results of the subsequent training of the pretrained Decision Transformer using the evolution strategy are shown in Figure 2. Even though our approach is capable of training the model, the progression (the improvement) is visibly slower than in the previous section without the pretraining – except for the initial phase.
Again, we explored the possibility of utilizing a smaller population for this training. As can be seen in Figure 3(b), this time, although the reliability of the training is again diminished, most of the time the algorithm works surprisingly well compared to how it works with the full population. Still, there was a single run in which the model did not manage to improve.
Next, we tried using standard (i.e. the same as in the previous experiment without the pretraining) values for the learning rate and noise deviation hyperparameters both with and without the VBN, as illustrated in Figure 3(b). Without the VBN the training showed no progression at all, and with the VBN the training started to show similar patterns towards the end of the training in a few cases as the one from Section III-B, yet still was maybe a bit slower.
(a) Evaluation results
(b) Wall-clock time
Figure 4: Median and quartiles of fitness values and wall-clock time for OpenAI-ES used on a Decision Transformer model for Atari game Hero in ten runs of the experiment. Figure 4(a) shows us the best-yet evaluation results after every iteration of the algorithm. In Figure 4(b), we can see the wall-clock time the training needed each iteration in seconds.
@@ III-D Decision Transformer - Atari
As a final experiment, we tested the ability of OpenAI-ES to train an even larger Decision Transformer in Atari games environment, more specifically, in the game Hero. Again, since the model is approximately three times bigger than in the previous experiments, we further doubled the population size for the algorithm. Even though Figure 4(a) shows that a gradual improvement of the model really does occur, Figure 4(b) reveals that it is accompanied by quite a significant increase in the time required to process a single iteration. This is due to a slower model inference, as the model is substantially larger, as well as longer episodes in the environment. (For the other experiments, the wall-clock time is not as interesting but can nonetheless be found alongside all the remaining collected data in the GitHub repository.) We did not include a gradient baseline for this experiment, as it primarily serves to showcase the much longer training durations required for such large models.
@@ IV Discussion
As we can see by comparing the experiments in Figure 2, the larger Decision Transformer model seems to have a fitness advantage during the training over the smaller feedforward model. But what is more important, OpenAI-ES appears to be capable of training it without any substantial problems, although the process takes longer. However, the training is less robust – during one experiment run, the model failed to achieve any significant improvement. This might have been caused simply by an inconvenient seed, when the perturbations tested during the training were not suitable. Or this might be caused by one design choice of the algorithm, where for better runtime the values for perturbations are pregenerated at the beginning of the training and so for the larger models we might have to generate more values at the beginning. Or maybe we might simply need to further increase the population size for greater robustness.
@@
Regarding our specific usage of the algorithm on Decision Transformers, we need to settle one thing. As we introduced in Section II-B, Decision Transformer operates with return-to-go tokens, which in the original paper [10] serve as a desired return the agent should achieve in a remainder of an episode. This works well when training in the original offline reinforcement learning manner, but when using the evolution strategy to train the Decision Transformers in our online setting, there is no pressure to take these tokens into account, and thus the tokens are then ignored, as can be seen in Table I. Nonetheless, we believe that only a minor modification of the algorithm may induce the agent to pay attention to these tokens. Let us propose an outline of such modifications for a possible future work. Each evaluation would consist of several subevaluations; each subevaluation would be assigned a desired return sampled from a $\mathcal{N}(\mu,\sigma)$, a normal distribution centered at $\mu$ with variance $\sigma$. The value of $\mu$ should be computed from the returns obtained during the previous iteration in such a manner that it is within an interval between the best return obtained and the mean return obtained during the last iteration, so an improvement is gradually made. The variance $\sigma$ would be a hyperparameter, or it could possibly be a variation of the returns obtained during the last iteration. The fitness of an individual would not be the mean of its returns obtained, as is the case in OpenAI-ES, but rather a decreasing function of the absolute values of the differences between the return obtained and the desired return in each subevaluation; ergo, the larger the differences, the smaller the fitness. This might be, e.g., the negative of a weighted sum of the differences.
@@
Even when doing so, we can see in Figure 2 that the model first gets “broken” – its performance worsens during first few iterations, the best-yet found does not improve – before it starts to improve again. Nevertheless, the improvement comes in terms of iterations much sooner than when no pretraing is utilized (and it somewhat works even with smaller population). Furthermore, we can observe by inspecting rollouts of both the initial and final models that those two show similar gaits. This similarity gets even more accentuated, when compared to the diversity in final behaviors that are yielded from the experiment in Section III-B, so when the algorithm is run without the pretraining. Therefore, we can clearly see that when initialized by the pretrained model the training clearly builds upon the seeded model, despite its initial deterioration.
@@
We hypothesize the reason for the above-mentioned deterioration at the beginning of the training from pretrained model, as well as the need for the aforementioned change of hyperparameters is the following: Gradient training – of which the behavior cloning is a representant – follows a different overall goal than the evolution strategy. It strives for the best possible parameters of the model it trains with regard to its loss function. In our case of behavior cloning, this means that we want to get our model to behave exactly as the one we are cloning towards. Nonetheless, this tells us nothing about how the model behaves when we perform a slight perturbation of parameters. Hence, after the perturbation, the model is likely to no longer function properly – or at least as well – anymore. So we can say that the model is brittle in a sense that even a slight change of parameters is likely to somehow break the model. This is, however, in stark contrast to how the evolution strategy operates and what it tries to achieve. The goal of the evolution strategy is not only that its current model performs as well as possible, but that we get the best possible behavior in expectation when sampling model parameters from its current distribution. Thus, we believe the models trained by the evolution strategy might be considered more robust – that might help, e.g., in areas where we train on precise hardware, but the model needs to be afterwards deployed on a cheaper hardware with less precision, where the weights of the model need to be rounded. Ergo, the algorithm needs to first, let us say, “robustify” the model and only then it can train it towards a better performance. This would be supported by the findings of Lehman et al. [29].
@@
Regardless, in the end, the results shown in Section III-B suggest that the pretraining is not really necessary. But then what explains why the problem stated at the beginning of Section III-C does not occur in practice? We think a possible answer is that the algorithm improves not only by shifting model weights towards the better performing individuals of the population, but when there are no really better individuals in the population, it does so mostly by shifting the weights away from the worst performing individuals, as the majority of the population consists of worse individuals. This might indicate why the algorithm does not really struggle with training larger and more complex models, where it is even harder to stumble onto a working solution just by adding some noise to the model parameters.
Finally, let us just briefly comment on Section III-D. The OpenAI-ES is capable of training even such larger models, but the larger the model the more computing power is required for the training to be concluded in a timely manner. Compared to sizes of transformer models powering current chatbots, the Decision Transformer used to learn the Atari game is miniscule, and even so it required a lot of wall-clock time to undertake just tens of iterations of the training on lower hundereds of CPUs. Hence, training really large transformer would require a gargantuan number of CPUs – but again, that is, at least in theory, not something unachievable.
@@ V Conclusion
We have investigated the ability of OpenAI-ES (and, in extension, of similar evolution strategies) to train larger and more complex models than the standard feedforward ones used in the literature so far in the reinforcement learning. We have also suggested a method to help this training by pretraining the model using behavior cloning towards some smaller, possibly weaker, yet easier to train model. However, the evolution strategy proved to be capable of training large models, whereas the pretraining showed multiple flaws during the experiments. Still, it helped us shed light onto the field of hybridizing gradients and evolution strategies, as we have discussed based on our experiments. It helped us understand why the sequential combination of first training using gradients and then proceeding to train using similar distribution-based evolution strategies might not work well. And it also allowed us to hypothesize that the distributional evolution strategies like OpenAI-ES may yield robust models, which are less prone, e.g., to rounding the weights due to low-precision hardware they might be deployed on.